# Nonsmooth Implicit Differentiation: Deterministic and Stochastic Convergence Rates

### v1

@@ E.2 Stochastic Case
$\qquad+4C^{2}_{E,\lambda}\big{(}1+\kappa\lVert D_{\Phi,2}(w(\lambda),\lambda)\rVert_{\sup}\big{)}^{2}\rho_{\lambda}(t)+\frac{1}{J_{1}}\big{(}\sigma_{E,2}+\sigma_{E,1}\kappa^{2}\lVert D_{\Phi,2}(w(\lambda),\lambda)\rVert^{2}_{\sup}\big{)}$
@@
The first part of the statement follows by noting that for NSID we have $\sigma_{\lambda}(k)=O(\kappa^{5}/k)$, where $\kappa=1/(1-q)$. The second part is immediate. ∎
@@ F.3 Details for the AID and ITD Experiments
We construct the synthetic dataset by sampling each element of the matrix $X\in\mathbb{R}^{n\times p}$ and the vector $w$ from a normal distribution. Subsequently, we set the non-informative features of $w$ to zero and we compute the vector $y$ as $y=Xw+\epsilon$, where $\epsilon_{i}$ is Gaussian noise with mean $0.1$ and unit variance. For this experiment we set $n=500$ and $p=100$ of which $30$ are informative.
@@ F.4 Details for the Stochastic Experiments
For the data poisoning setup use the MNIST dataset. We split the MNIST original train set into $30K$ example for training and $30K$ examples for validation. Additionally, we perform a random split of the training set into $X\in\mathbb{R}^{n\times p}$ and $\tilde{X}\in\mathbb{R}^{n^{\prime}\times p}$, with $p=784$ representing the number of features for MNIST images. Notably, $n^{\prime}=9K$ denotes the number of corrupted examples. It is essential to highlight that $\Gamma\in\mathbb{R}^{n^{\prime}\times p}$ and $n^{\prime}p$ is approximately $7$ million, posing a significant challenge for derivative estimation using zero-order methods. For data poisoning, we observed that the theoretical value for the step size was too conservative and hence we multiply it by $10$, to have improved convergence. We set the regularization parameters $\lambda=(0.02,0.1)$ since with this setup, the final uncorrputed linear model achieves a validation accuracy of around $80\%$ with around $90\%$ of components set to zero. We note that NSID and SID require to choose the step sizes $(\eta_{i})$, which we found to be difficult, since the theoretical values are often conservative estimates for this problem. We try two policies: constant and decreasing (as $\Theta(1/i)$) step sizes, indicated with “const” and “dec” after the method name respectively. Note that only when the step sizes are decreasing NSID is guaranteed to converge. To simplify the setup, we always set them equal at $i=0$. Moreover, we set the step size of NSID equal to that of SID, when they use the same step sizes policy.

### v2

@@ E.2 Stochastic Case
$\qquad+3(1+\kappa\lVert D_{\Phi,2}(w(\lambda),\lambda)\rVert_{\sup})^{2}\Big{(}C^{2}_{E,\lambda}\Delta^{2}_{t}+\frac{\sigma_{E,1}+\sigma_{E,2}}{J_{1}}\Big{)}\big{)}$
@@
The first part of the statement follows by noting that for NSID we have $\sigma_{\lambda}(k)=O(\kappa^{5}/k)$, where $\kappa=1/(1-q)$. The second and last result are immediate. ∎
@@ F.3 Details for the AID and ITD Experiments
We construct the synthetic dataset by sampling each element of the matrix $X\in\mathbb{R}^{n\times d}$ and the vector $w$ from a normal distribution. Subsequently, we set the non-informative features of $w$ to zero and we compute the vector $y$ as $y=Xw+\epsilon$, where $\epsilon_{i}$ is Gaussian noise with mean $0.1$ and unit variance. For this experiment we set $n=500$ and $p=100$ of which $30$ are informative.
@@ F.4 Details for the Stochastic Experiments
For the data poisoning setup we use the MNIST dataset. We split the MNIST original train set into $30K$ example for training and $30K$ examples for validation. Additionally, we perform a random split of the training set into $X\in\mathbb{R}^{n\times d}$ and $\tilde{X}\in\mathbb{R}^{n^{\prime}\times p}$, with $p=784$ representing the number of features for MNIST images. Notably, $n^{\prime}=9K$ denotes the number of corrupted examples. It is essential to highlight that $\Gamma\in\mathbb{R}^{n^{\prime}\times p}$ and $n^{\prime}p$ is approximately $7$ million, posing a significant challenge for derivative estimation using zero-order methods. For data poisoning, we observed that the theoretical value for the step size was too conservative and hence we multiply it by $10$, to have improved convergence. We set the regularization parameters $\lambda=(0.02,0.1)$ since with this setup, the final uncorrputed linear model achieves a validation accuracy of around $80\%$ with around $90\%$ of components set to zero. We note that NSID and SID require to choose the step sizes $(\eta_{i})$, which we found to be difficult, since the theoretical values are often conservative estimates for this problem. We try two policies: constant and decreasing (as $\Theta(1/i)$) step sizes, indicated with “const” and “dec” after the method name respectively. Note that only when the step sizes are decreasing NSID is guaranteed to converge. To simplify the setup, we always set them equal at $i=0$. Moreover, we set the step size of SID equal to that of NSID, when they use the same step sizes policy.